OpenStudy (anonymous):

1.The surface area of a rectangular prism is 150 cm². If all the sides are tripled, what will the new surface area be?

OpenStudy (anonymous):

Surface Area of a Rectangular Prism = 2ab + 2bc + 2ac if If all the sides are tripled, 3a,3b,3c new surface area be 2(3a)(3b)+2(3b)(3c)+2(3c)(3a)=9(2ab + 2bc + 2ac) 9 times of prevois area

OpenStudy (anonymous):

understtod?

OpenStudy (anonymous):

so would it be 9*150=1,350

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnk you

OpenStudy (anonymous):

giv me medal

OpenStudy (anonymous):

giv me medal if u lik my answer

OpenStudy (anonymous):

how do u do tht again i forgot

OpenStudy (anonymous):

click that blue button

OpenStudy (anonymous):

ok